OpenStudy (anonymous):

Find the limit of the following sequences or determine that the limit does not exist. {Bn} where Bn = {n/(n+1) if n < or equal to 5000 {ne^-n if n>5000

OpenStudy (anonymous):

Since it is a sequence, we're taking the limit as \(n\to\infty\).

OpenStudy (anonymous):

yes ok

OpenStudy (anonymous):

We're going to use that \(ne^{-n}\) case since that is the case used as we approach \(\infty\).

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

Looking at it straight up, we get \(\infty e^{-\infty}\to \infty\times 0\) which is an indeterminate form.

OpenStudy (anonymous):

Well \[ \frac{n}{e^n} \]Can be solved with l'Hospital's rule.

OpenStudy (anonymous):

idk the d/dx of e^n

OpenStudy (anonymous):

Suppose \[ f(n)=a_n\]where \(f(n)\) is a continuous function. Then I'm pretty sure: \[ \lim_{n\to \infty }a_n = \lim_{x\to \infty} f(x) \]when the limit exists.

OpenStudy (anonymous):

In fact, the definition of a limit and sequence are the same so yeah, they are the same.

OpenStudy (anonymous):

So treat it like a continuous function and differentiate with respect to \(n\) .

OpenStudy (anonymous):

whats the final answer?

OpenStudy (anonymous):

What is the derivative?

OpenStudy (anonymous):

1/e^x.

OpenStudy (anonymous):

Okay so what does that limit approach?

OpenStudy (anonymous):

5000